OpenStudy (anonymous):

The pH of a juice drink is 2.6 Find the concentration of hydrogen ions in the drink. a.2.6 b.2.5*10^-3 c.10^-2.6 d.2.5*10^-3

OpenStudy (australopithecus):

do you have the formula?

OpenStudy (anonymous):

No I don't . :/

OpenStudy (australopithecus):

pH = -log_10[H+]

OpenStudy (australopithecus):

I should really have that formula memorized oh well

OpenStudy (australopithecus):

that is the formula, do you know how to solve it now?

OpenStudy (australopithecus):

http://www.ausetute.com.au/phscale.html

OpenStudy (australopithecus):

I can show you how to solve it if the answer is no

OpenStudy (australopithecus):

Also note pH = -log[H+] as it is log base 10

OpenStudy (anonymous):

Still confused and the answer is no .

OpenStudy (australopithecus):

alright note that log(x) = y is the same as x = 10^(y)

OpenStudy (anonymous):

Okay

OpenStudy (australopithecus):

Also note that 10^(log(x)) = x 7^(log_7(x)) etc.. pH = -log_10[H+] So we know the pH is equal to pH = 2.6 thus 2.6 = -log_10[H+] multiply both sides by -1 -1 * 2.6 = -log_10[H+] * -1 -2.6 = log_10[H+] Now make both sides of the equation exponents to the base 10 to eliminate the logarithm 10^(-2.6) = 10^(log_10[H+]) 10^(-2.6) = [+H] REMEMBER SQUARE BRACKETS MEAN CONCENTRATION (probably Molarity) so we know the concentration of +H is 10^(-2.6) M

OpenStudy (australopithecus):

I hope this helps you, logarithms are pretty important to understand I think at least for solving some basic chemistry formulas

OpenStudy (anonymous):

Chemistry formulas are important to understand but I got this question on Algebra 2 HW

OpenStudy (australopithecus):

Weird to get a question with a specific formula like this, although I'm sure it makes sense if you analyze the question. Logarithmic rules 1) log_b(mn) = log_b(m) + logb(n) 2) log_b(m/n) = log_b(m) – logb(n) 3) log_b(m^(n)) = n · log_b(m) 4) log_3(3) = 1 5) log_e(x) = ln(x) 6) log(x) = y is the same as x =10^(y) 7) log_10(x) = log(x) 8) log_e(x) = ln(x) 9) e^(ln(x)) = x

OpenStudy (australopithecus):

here are the rules for logarithms hope they help you

OpenStudy (australopithecus):

do you understand how to solve it now?

OpenStudy (australopithecus):

Note that 10^(-2.6) = 0.0025M which is a pretty high concentration of H+ ions which is expected of a solution that has a pH that low. So this answer is reasonable

OpenStudy (anonymous):

So 4 log x - 6 log (x+2) is an example of rule #2 ?

OpenStudy (anonymous):

Ya I got that on my TI .

OpenStudy (australopithecus):

log_b(m/n) = log_b(m) – logb(n) it is an example if the original logarithm was log(x^(4)/(x+2)^(6))

OpenStudy (australopithecus):

as you can simplify it down to what you have written there

OpenStudy (anonymous):

But I have to simplify it to a single logarithm .

OpenStudy (australopithecus):

so just follow my rules start with the logarithms individually

OpenStudy (australopithecus):

or the rules they aren't mine :)

OpenStudy (australopithecus):

sorry I made a mistake on rule 2 might have made it confusing 2) log_b(m/n) = log_b(m) – log_b(n)

OpenStudy (australopithecus):

4 log x - 6 log (x+2) for the first one see that log(x^(4)) = 4log(x) and log((x+2)^(6)) = 6log(x+2)

OpenStudy (australopithecus):

Notice the negative in front of the logarithm indicates that the exponent is negative - 6 log (x+2) so - 6 log (x+2) = log((x+2)^(-6)) This is significant because x^(-1) = 1/x x^(-56) = 1/x^(56) 1/x^(-1) = x/1 etc so we know right away when we combine these logarithms that it is going to be in the denominator

OpenStudy (australopithecus):

The last comment is just for interest you can blindly apply the rules but understanding why things are the way they are is always nice

OpenStudy (anonymous):

I put it into my calculatorr and got -2.5

OpenStudy (australopithecus):

what?

OpenStudy (anonymous):

I put that problem into my calculator and came out with -2.5